OpenStudy (zyberg):

Need some help with mathematical logic: Prove that (A → B) ~ (¬A ∨ ¬B). My book says that it is always true, however i can not understand why. I tried making tables of truth, but it is not getting anywhere...

OpenStudy (zyberg):

My initial table of truth: A B "A" "B" ("A" ∨ "B") (A → B) 1 T T F F F T 2 T F F T T F 3 F T T F T T 4 F F T T T T So, (A → B) ~ ("A" ∨ "B"): 1 F 2 F 3 T 4 T

OpenStudy (zyberg):

Where did I make a mistake?

OpenStudy (loser66):

I don't understand your problem. Can you take a snapshot?

OpenStudy (zyberg):

@Loser66 I had found a way to denote the negation (that's what I had meant with inverse). Now question is exactly as what was stated.

OpenStudy (zyberg):

@sleepyjess

OpenStudy (zyberg):

I have rewritten the question and my attempt to solve in a fancier manner (with table and such) on MSE: http://math.stackexchange.com/questions/1925326/mathematical-logic-problem-about-proof

ganeshie8 (ganeshie8):

Your textbook is wrong

ganeshie8 (ganeshie8):

Below is correct : (A → B) ~ (¬A ∨ B)

OpenStudy (zyberg):

@ganeshie8 I think it is the first time I had spotted a mistake in my textbooks, but it's awesome to know that I wasn't wrong (tried doing the same thing with (A → B) ~ (¬A ∨ B) and I got the right result) Thank you very much for spotting it! :)

ganeshie8 (ganeshie8):

Nice, looking at truth table may not satisfy you. You should try and get convinced of the result in other ways..

OpenStudy (zyberg):

@ganeshie8 what do you mean?

ganeshie8 (ganeshie8):

\(A \to B\) is false only when A is true and B is false, yes ?

ganeshie8 (ganeshie8):

If it is raining, you will carry umbrella. A : it is raining B : you carry umbrella

ganeshie8 (ganeshie8):

Notice that the statement says what you do when it rains. (A is true) It doesn't say anything about your actions when it doesn't rain. (A is false)

ganeshie8 (ganeshie8):

That means the statement holds true no matter what you do when it doesn't rain.

OpenStudy (zyberg):

That's quite an interesting example! Thank you for providing it! When I first looked at the truth table it looked like I was supposed to memorize it (as always - a sign that I am not understanding something), but after some googling I understood that it is quite hm... logical. It is really fascinating how mathematical logic is connected so much with every day life (it sounds quite stupid, when put into words, but I mean that it's quite interesting how we can put everything in a few tables and it explains a lot of things that happen every day). It was very interesting for me to learn that we can apply a few mathematical rules to simplify electrical circuits! :) Sorry to bother you, @ganeshie8 but could you think of a way to represent "~" with some kind of real life example (like the one with umbrella)?

ganeshie8 (ganeshie8):

By "~", do you mean "negation" or a "tautology" ?

OpenStudy (zyberg):

I think that I am talking about tautology.

ganeshie8 (ganeshie8):

tautology is simply another way to say the same statement

ganeshie8 (ganeshie8):

Consider below statements statement1 : I am your sibling statement2 : You are my sibling

ganeshie8 (ganeshie8):

Would you agree that they are logically equivalent ?

OpenStudy (zyberg):

Yes.

ganeshie8 (ganeshie8):

We call it a tautology and represent it as : statement1 \(\iff\) statement2

ganeshie8 (ganeshie8):

Proper symbol to use is \(\iff\) May I know why you're using ~ for tautology ?

OpenStudy (zyberg):

Oh! You were right saying that truth tables are not the right way to look into this! Well, my book tells me that it is either double arrow (that you used) or ~, but I have no idea how to write double arrow, so I am using ~.

ganeshie8 (ganeshie8):

Okay, I see... good. So a tautology is when you have two logically equivalent statements. You may think of that sibling example and double implication when you see a tautology

ganeshie8 (ganeshie8):

Do you see why below is a tautology ? (A → B) ~ (¬A ∨ B)

ganeshie8 (ganeshie8):

In other words, why the expression to the left of ~ is logically equivalent to the expression on the right ?

OpenStudy (zyberg):

Yes, both (A → B) and (¬A ∨ B) produce the same results, so they are logically always equivalent.

ganeshie8 (ganeshie8):

Can you see that with out using truth tables ?

ganeshie8 (ganeshie8):

familiar with Demorgan laws ?

OpenStudy (zyberg):

No, I can't. And I am not familiar with it. Truth to be told, it was the very first problem I had tried to solve using mathematical logic and all those symbols.

ganeshie8 (ganeshie8):

Demorgan Law1 : \(\neg (A \lor B)\) is logically equivalent to \(\neg A \land \neg B\)

ganeshie8 (ganeshie8):

When you're free, may be cookup few examples and try to figure out how/why demorgan law works..

ganeshie8 (ganeshie8):

Do not depend on truth tables so much, use and/or/not logic instead

ganeshie8 (ganeshie8):

Truth tables prove things, but they don't help you learn anything new..

OpenStudy (zyberg):

What do you mean by and/or/not logic?

ganeshie8 (ganeshie8):

Every statement in logic can be represented using those 3 symbols

ganeshie8 (ganeshie8):

Let me show you quick how (A → B) and (¬A ∨ B) are equivalent using simple words

OpenStudy (zyberg):

I am very eager to see that :) (I had worked it out using tables again)

ganeshie8 (ganeshie8):

We know that A → B is false only when A is true and B is false, yes ?

ganeshie8 (ganeshie8):

its raining outside, but you're not carrying umbrella

ganeshie8 (ganeshie8):

That is the only way to violate the statement : If it is raining, you will carry umbrella.

OpenStudy (zyberg):

Yes, I understand that.

ganeshie8 (ganeshie8):

We know that A → B is false only when A is true and B is false. That is same as saying A → B is always true except when A is true and B is false.

ganeshie8 (ganeshie8):

Next look at "or" logic. "X or Y" is always true except when X is false and Y is false, yes ?

OpenStudy (zyberg):

Yes.

OpenStudy (zyberg):

(Sorry for delays before my answers - it's very unusual for me to see everything put into words. I played around with and, or, not logic with programming, but never tried to put everything into words and it stayed as an idea without a clear formulation in my head)

OpenStudy (zyberg):

OH! Now I see what you meant. It's quite interesting! I am going to try and solve all the problems using such method without really relying on the tables (although table is a description of and/or/not, so in a way it is still relying on it ;) ). Thanks for showing it to me!

OpenStudy (zyberg):

Basically, what you meant is to look at the cases when it's false, yes?

ganeshie8 (ganeshie8):

Yes, there is only one false case. Looking at 1 case is easier than looking at 3 cases, right ?

ganeshie8 (ganeshie8):

Saying "last week, it didn't rain only on sunday" is easier than saying "last week, it rained only on monday, tuesday,..., saturday"

ganeshie8 (ganeshie8):

both convey the same information, but the first statement certainly looks simpler

OpenStudy (zyberg):

How would you go about simplifying the logical statements, @ganeshie8? For example this one: ((A1 v A3) ^ A2) v (A1 ^ ¬A2 ^ A3) v (¬A1 ^ ¬A2 ^ A3). Is it possible to simplify this without a need of truth tables?

ganeshie8 (ganeshie8):

Yes, that is specially made to simplify with out using truth tables.

ganeshie8 (ganeshie8):

First lets look at few things quick

ganeshie8 (ganeshie8):

1) what's the truth value of the expression \(A \lor \neg A\) ? Does it evaluate to true or false ?

ganeshie8 (ganeshie8):

example : I'll be happy if it rains or doesn't rain tomorrow. \(A\) : it rains \(\neg A\) : doesn't rain Would I be happy tomorrow in regards to raining/not raining ?

OpenStudy (zyberg):

@ganeshie8, sorry that I hadn't responded - was in school. it would always be True.

ganeshie8 (ganeshie8):

Let's use 1 to mean false and 0 to mean true

ganeshie8 (ganeshie8):

With above notation, \(A \lor \neg A = 1\)

ganeshie8 (ganeshie8):

Let me ask you another question. In algebra, what do below expressions refer to ? 1) xy 2) x + y

OpenStudy (zyberg):

Well, it is multiplication and addition, isn't it, @ganeshie8 ?

ganeshie8 (ganeshie8):

Yes, why are we using a special symbol "+" for addition and not using any symbol for multiplication ?

ganeshie8 (ganeshie8):

Because it looks simple and clear if we don't use \(*\) or \(\times \) to mean multiplication

ganeshie8 (ganeshie8):

We can do the same in logic. It may look weird in the starting, but after using it 2-3 times, you will get used to it.

ganeshie8 (ganeshie8):

Lets use "+" to refer to "or" logic and "nothing" to refer to "and" logic

ganeshie8 (ganeshie8):

For example : \(A \lor B \) in our new compact notation becomes \(A+B\) \(A \land B \) in our new compact notation becomes \(AB\)

OpenStudy (zyberg):

Oh my! So, the problem gets down to: ((A1 + A3)A2) + (A1¬A2A3) + (¬A1 ¬A2A3), right?

OpenStudy (zyberg):

It actually looks a lot less scary with arithmetic notation, hehe.

ganeshie8 (ganeshie8):

Looks nice, isn't it ? One more thing before getting back to simplifying it

ganeshie8 (ganeshie8):

Let's use \('\) to refer to the negation \(\neg \)

ganeshie8 (ganeshie8):

For example \(\neg A\) in our compact notation would be \(A'\)

OpenStudy (zyberg):

Okay, like in physics.

ganeshie8 (ganeshie8):

Btw, this notation is not invented by me. All engineers use this notation..

ganeshie8 (ganeshie8):

Engineers like to simplify things, so trust that this new notation is going to help you understand logic faster and better ;)

OpenStudy (zyberg):

Well, it already looks easier ;) (A1 + A3)A2 + A1A2`A3 + A1`A2`A3 = (A1 + A3)A2 + A2`A3(A1 + A1`), as I see.

ganeshie8 (ganeshie8):

Perfect!

OpenStudy (zyberg):

Since x + x` = 1, like we talked earlier we can just "erase" that A1 + A1`?

ganeshie8 (ganeshie8):

Yes, replace that by "1"

ganeshie8 (ganeshie8):

Don't forget that \(AB\) means \(A \land B\)

OpenStudy (zyberg):

(A1 + A3)A2 + A2`A3 = A1A2 + A2A3 + A2`A3 = A1A2 + A3(A2 + A2`), so the same thing and we get A1A2 + A3 and in symbolic notation it would be A1 ^ A2 v A3, it's same as the answer in the book (though, there are brackets around A1 and A2, how did they appear here?).

OpenStudy (zyberg):

Seems that it's a lot easier to use the engineering notation :) Maybe you know some more examples of such simplifications? It seems that my book provides only "prove X" type of problems and I am still not very confident with this simplification thing.

ganeshie8 (ganeshie8):

Yes, just so you know, the entire semiconductor industry is based on this simplification. We use computers to simplify these boolean expressions invonvling millions of or/and/not gates. It gets messy to simplify manually once the number of variables in the expression become more than 4..

ganeshie8 (ganeshie8):

Let me cook up an interesting expression to simplify..

ganeshie8 (ganeshie8):

xz' + xyz' = ?

OpenStudy (zyberg):

hm, xz`(1 + y), since + is or, meaning that it is always true? since, we are comparing true and something with or statement. so, we get 1 + y = 1, then xz` is left, so x ^ z`. Was that right?

ganeshie8 (ganeshie8):

Excellent! one more ?

OpenStudy (zyberg):

Yes, I would be happy to see one more.

ganeshie8 (ganeshie8):

x + x'y = ?

OpenStudy (zyberg):

Mmm... Perhaps we can somehow make x to be x` plus or times something? Since we don't(?) have minus it would make sense, but I can't really figure out what to exchange x for...

ganeshie8 (ganeshie8):

Ahh sorry, it's a tricky one. Let me rephrase the q : Prove x + x'y = x + y

ganeshie8 (ganeshie8):

use any method

ganeshie8 (ganeshie8):

You may also argue in simple sentences and convey your understanding. Its considered a valid proof if it follows valid reasoning...

OpenStudy (zyberg):

I see that it's not possible to add things to both sides, to prove this. I also see a weird thing: 1 + x = 1, then I should be able to multiply anything by 1 + x without changing a thing, however: x + x`y(1 + x) = x + y x + x`y + y = x + y Meaning that x`y = 0, which isn't really true? (It's weird to me since I arrived at this conclusion a few times, but it seems to be false to me)

ganeshie8 (ganeshie8):

You're wrongly assuming that x'x = 1

ganeshie8 (ganeshie8):

Check the red parts, they are not same. x + \(\color{red}{x`y(1 + x)}\) = x + y x + \(\color{red}{x`y + y}\) = x + y